# Circular Future of Yukawa Unification in Supersymmetric Pati-Salam Model: Stop on Neutralino

### 26 Mar 2026 v1

@@ 3 Scanning Procedure, Experimental Constraints and Detector Analyses
In the analyses for the collider implications, we employ MadGraph5_aMC@NLO (v3.6.3) Alwall:2014hca to generate the signal and relevant background events at the parton level. Following the analyses in Ref.CMS:2025ttk , we set the five-flavor-scheme (5FS) and use NNPDF23LO1 NNPDF:2017mvq at this step. The generated events are transferred to Pythia 8.3 Sjostrand:2014zea to complete the final states by parton showering, hadronization and decaying unstable particles. The detector response is also included in our simulations with the use of Delphes-3.5.0 deFavereau:2013fsa . Since we expect heavy stops which are more likely beyond the reach of the current experiments, we simulate similar analyses reported in Ref.CMS:2025ttk for FCC-hh with 100 TeV Center of Mass (COM) energy, by employing FCChh_II.tcl card within Delphes, which is generated suitably with the official design of FCC-hh. We discuss our results at this step in terms of the statistical significance for the signal events, which is calculated as follows Cowan:2010js ; Kumar:2015tna :
@@ 5.2 Stop On Neutralino
Before proceeding into detailed detector analyses, we present our results for mass scales for stop in Figure 9 in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ and $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ planes, where $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})\equiv m_{\tilde{t}_{1}}-m_{\tilde{\chi}_{1}^{0}}$. The color coding is the same as in the top planes of Figure 6. The curves represent the exclusion from the experimental analyses CMS:2021beq ; CMS:2025ttk . The YU solutions (orange) around the exclusion curve can yield stops as light as about 1.5 TeV which are slightly beyond the exclusion limit as seen in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ plane. The studies have projected HL-LHC sensitivity to probe such stop solutions up to about 2.2 TeV Baer:2023uwo ; Baer:2025zqt , and the FCC-hh experiments will more likely improve the results up to about 3 TeV Altin:2020qmu .
The interesting observation in these results can be realized in the region where the stop and LSP neutralino are nearly degenerate in mass. Such solutions are not present in the holomorphic case Hussain:2025bxp . Even though such solutions can escape from the sensitivity of the analyses mentioned above, they can be subjected to the analyses over compressed spectra in which the stop does not weigh much heavier than LSP neutralino. Recently revealed results from these analyses CMS:2025ttk are represented with the dashed and solid curves in the $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ plane. However, these analyses can probe the stop up to about 700 GeV yet, while the YU solutions appear in the region with $m_{\tilde{t}_{1}}\gtrsim 1.4$ TeV. These solutions are quite beyond the sensitivities of the current collider experiments, but they can be tested in future experiments, which are proposed to operate with much higher center of mass energies such as FCC-hh.
@@
Figure 10 displays the significance for the events diagrammatically represented in Figure 2. The significance calculated as given in Eq.(10) projected to the FCC-hh experiments of 100 TeV COM. The color coding is the same as in the top planes of Figure 6. The horizontal dashed lines represent the significance equal to 1, 2 and 3 from bottom to top, respectively. The left plane shows the significance for $\mathcal{L}=10$ fb-1. The YU solutions (orange) can yield stop-pair production events with $\mathcal{Z}\simeq 1$, which can be probed in FCC experiments up to about $68\%$ CL. If one also requires these solutions to be compatible with the Planck measurements on relic abundance of LSP neutralino (red), the maximum significance can be realized at about 0.5. The right panel shows that these results can slightly be improved when the integrated luminosity reaches to about 139 fb-1 as $\mathcal{Z}\simeq 1.2$ for the YU solutions (orange) and $\mathcal{Z}\simeq 0.8$ when the YU solutions are consistent with the DM observations (red).
@@ 5.3 FCC Probe of Compressed Stop-Neutralino
Based on the results discussed with the plots in Figure 11 we impose the cuts on the kinematic variables subsequently as listed in Table 3. The final results are obtained after the fourth set of the cut flow, but we list them as different sets to discuss the effects from each cut in details. Recall that these cuts are imposed after the pre-selection rules are applied. The calculated number of events for FCC-hh of 100 TeV COM are listed in Table 4 after each set. The effectiveness of each cut set on the signal processes is parametrized by $\epsilon_{S}$, which is simply the ratio of number of signal events as
@@
Table 4: The number of events for the signal and background processes realized at FCC-hh experiments of 100 TeV COM and $100$ fb-1 integrated luminosity together with the effectiveness of cut flows and resultant significance for the signal processes.
It can be seen easily from the results given in Table 4 that the cut on the missing energy immediately increases the significance of the signal events by about 10 times, while the other sets can also provide minor improvements. Eventually, only about $16\%$ of the signal events can survive after cuts in Set 4, which corresponds to $\mathcal{Z}\simeq 5.2$. Recall that we assume $\mathcal{L}=100$ fb-1 in calculation of significance, which can be collected on the first days of FCC-hh operation. In this context, one can expect significantly improved results as the integrated luminosity increases, as shown in Figure 12. According to the correlation between the significance and integrated luminosity, the compressed spectra of the stop and LSP neutralino can be probed up to about $3\sigma$ when $\mathcal{L}\simeq 32$ fb-1, and $5\sigma$ probe can be realized for $\mathcal{L}\simeq 97$ fb-1.
@@
Even though the cut-based analyses can yield a strong probe at the FCC-hh experiments, the results obtained above can be improved further through Boosted Decision Tree (BDT) analyses. We employ the BDT algorithm based on XGBoost (eXtreme Gradient Boost) trained with $10^{5}$ events for each of the signal and background processes generated independently. In addition to the variables described above, the events are transferred to the BDT scan with $m_{CT}$ Tovey:2008ui , $m_{T}\!\left(\ell,E_{T}^{\mathrm{miss}}\right)$ ATLAS:2024cmj , $m_{T2}^{W}$ Bai:2012gs , $m(b,l)$, $m_{T}\left(b_{i},E_{T}^{\mathrm{miss}}\right)$ ATLAS:2017drc . Note that the events are refined only by applying the pre-selection rules without any further cut on the variables.
@@
After training, BDT can identify the signal processes over the background quite efficiently as shown in the left panel of Figure 13. The best score in our analyses is obtained as $\tau=0.98$, which is tested over the signal significance displayed in the right plane. With the best score in BDT, we realize a significant improvement that the analyses can probe the signal up to about $3\sigma$ with the integrated luminosity of about $19.4$ fb-1, and $5\sigma$ when $\mathcal{L}\simeq 54.4$ fb-1. The significance of the signal in correlation with the integrated luminosity is given in Figure 14. One can expect from these results that the compressed spectra for the stop and neutralino are more likely to be probed on the first days when FCC-hh starts being operated.
@@ 6 Conclusion
In our analyses, we also require the relic density constraint to be satisfied, and we find the lightest possible stop weigh about 1.7 TeV of about 200 GeV mass difference with the LSP neutralino. Such solutions are clearly beyond the reach of the current analyses, but they can be tested in future experiments such as those proposed in FCC-hh with 100 TeV COM. We discuss possible probes for these solutions by considering the collider analyses projected into FCC-hh. In these analyses, we consider semi-leptonic final states emerging from stop pair-production. With suitable selections described with cuts on the kinematic variables, we realize statistical significance at about $5\sigma$ when the integrated luminosity is increased to about $100$ fb-1. We also perform a BDT analysis which results in such a large significance when the experimental analyses have lower luminosity as $\mathcal{L}\simeq 54$ fb-1. Such luminosity values can be reached on the first days of FCC-hh experiments.

### 30 Mar 2026 v2

@@ 3 Scanning Procedure, Experimental Constraints and Detector Analyses
In the analyses for the collider implications, we employ MadGraph5_aMC@NLO (v3.6.3) Alwall:2014hca to generate the signal and relevant background events at the parton level. Following the analyses in Ref.CMS:2025ttk , we set the five-flavor-scheme (5FS) and use NNPDF23LO1 NNPDF:2017mvq at this step. The generated events are transferred to Pythia 8.3 Sjostrand:2014zea to complete the final states by parton showering, hadronization and decaying unstable particles. The detector response is also included in our simulations with the use of Delphes-3.5.0 deFavereau:2013fsa . Since we expect heavy stops which are more likely beyond the reach of the current experiments, we simulate similar analyses reported in Ref.CMS:2025ttk for FCC-$h$ with 100 TeV Center of Mass (COM) energy, by employing FCChh_II.tcl card within Delphes, which is generated suitably with the official design of FCC-$h$. We discuss our results at this step in terms of the statistical significance for the signal events, which is calculated as follows Cowan:2010js ; Kumar:2015tna :
@@ 5.2 Stop On Neutralino
Before proceeding into detailed detector analyses, we present our results for mass scales for stop in Figure 9 in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ and $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ planes, where $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})\equiv m_{\tilde{t}_{1}}-m_{\tilde{\chi}_{1}^{0}}$. The color coding is the same as in the top planes of Figure 6. The curves represent the exclusion from the experimental analyses CMS:2021beq ; CMS:2025ttk . The YU solutions (orange) around the exclusion curve can yield stops as light as about 1.5 TeV which are slightly beyond the exclusion limit as seen in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ plane. The studies have projected HL-LHC sensitivity to probe such stop solutions up to about 2.2 TeV Baer:2023uwo ; Baer:2025zqt , and the FCC-$h$ experiments will more likely improve the results up to about 3 TeV Altin:2020qmu .
The interesting observation in these results can be realized in the region where the stop and LSP neutralino are nearly degenerate in mass. Such solutions are not present in the holomorphic case Hussain:2025bxp . Even though such solutions can escape from the sensitivity of the analyses mentioned above, they can be subjected to the analyses over compressed spectra in which the stop does not weigh much heavier than LSP neutralino. Recently revealed results from these analyses CMS:2025ttk are represented with the dashed and solid curves in the $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ plane. However, these analyses can probe the stop up to about 700 GeV yet, while the YU solutions appear in the region with $m_{\tilde{t}_{1}}\gtrsim 1.4$ TeV. These solutions are quite beyond the sensitivities of the current collider experiments, but they can be tested in future experiments, which are proposed to operate with much higher center of mass energies such as FCC-$h$.
@@
Figure 10 displays the significance for the events diagrammatically represented in Figure 2. The significance calculated as given in Eq.(10) projected to the FCC-$h$ experiments of 100 TeV COM. The color coding is the same as in the top planes of Figure 6. The horizontal dashed lines represent the significance equal to 1, 2 and 3 from bottom to top, respectively. The left plane shows the significance for $\mathcal{L}=10$ fb-1. The YU solutions (orange) can yield stop-pair production events with $\mathcal{Z}\simeq 1$, which can be probed in FCC experiments up to about $68\%$ CL. If one also requires these solutions to be compatible with the Planck measurements on relic abundance of LSP neutralino (red), the maximum significance can be realized at about 0.5. The right panel shows that these results can slightly be improved when the integrated luminosity reaches to about 139 fb-1 as $\mathcal{Z}\simeq 1.2$ for the YU solutions (orange) and $\mathcal{Z}\simeq 0.8$ when the YU solutions are consistent with the DM observations (red).
@@ 5.3 FCC Probe of Compressed Stop-Neutralino
Based on the results discussed with the plots in Figure 11 we impose the cuts on the kinematic variables subsequently as listed in Table 3. The final results are obtained after the fourth set of the cut flow, but we list them as different sets to discuss the effects from each cut in details. Recall that these cuts are imposed after the pre-selection rules are applied. The calculated number of events for FCC-$h$ of 100 TeV COM are listed in Table 4 after each set. The effectiveness of each cut set on the signal processes is parametrized by $\epsilon_{S}$, which is simply the ratio of number of signal events as
@@
Table 4: The number of events for the signal and background processes realized at FCC-$h$ experiments of 100 TeV COM and $100$ fb-1 integrated luminosity together with the effectiveness of cut flows and resultant significance for the signal processes.
It can be seen easily from the results given in Table 4 that the cut on the missing energy immediately increases the significance of the signal events by about 10 times, while the other sets can also provide minor improvements. Eventually, only about $16\%$ of the signal events can survive after cuts in Set 4, which corresponds to $\mathcal{Z}\simeq 5.2$. Recall that we assume $\mathcal{L}=100$ fb-1 in calculation of significance, which can be collected on the first days of FCC-$h$ operation. In this context, one can expect significantly improved results as the integrated luminosity increases, as shown in Figure 12. According to the correlation between the significance and integrated luminosity, the compressed spectra of the stop and LSP neutralino can be probed up to about $3\sigma$ when $\mathcal{L}\simeq 32$ fb-1, and $5\sigma$ probe can be realized for $\mathcal{L}\simeq 97$ fb-1.
@@
Even though the cut-based analyses can yield a strong probe at the FCC-$h$ experiments, the results obtained above can be improved further through Boosted Decision Tree (BDT) analyses. We employ the BDT algorithm based on XGBoost (eXtreme Gradient Boost) trained with $10^{5}$ events for each of the signal and background processes generated independently. In addition to the variables described above, the events are transferred to the BDT scan with $m_{CT}$ Tovey:2008ui , $m_{T}\!\left(\ell,E_{T}^{\mathrm{miss}}\right)$ ATLAS:2024cmj , $m_{T2}^{W}$ Bai:2012gs , $m(b,l)$, $m_{T}\left(b_{i},E_{T}^{\mathrm{miss}}\right)$ ATLAS:2017drc . Note that the events are refined only by applying the pre-selection rules without any further cut on the variables.
@@
After training, BDT can identify the signal processes over the background quite efficiently as shown in the left panel of Figure 13. The best score in our analyses is obtained as $\tau=0.98$, which is tested over the signal significance displayed in the right plane. With the best score in BDT, we realize a significant improvement that the analyses can probe the signal up to about $3\sigma$ with the integrated luminosity of about $19.4$ fb-1, and $5\sigma$ when $\mathcal{L}\simeq 54.4$ fb-1. The significance of the signal in correlation with the integrated luminosity is given in Figure 14. One can expect from these results that the compressed spectra for the stop and neutralino are more likely to be probed on the first days when FCC-$h$ starts being operated.
@@ 6 Conclusion
In our analyses, we also require the relic density constraint to be satisfied, and we find the lightest possible stop weigh about 1.7 TeV of about 200 GeV mass difference with the LSP neutralino. Such solutions are clearly beyond the reach of the current analyses, but they can be tested in future experiments such as those proposed in FCC-$h$ with 100 TeV COM. We discuss possible probes for these solutions by considering the collider analyses projected into FCC-$h$. In these analyses, we consider semi-leptonic final states emerging from stop pair-production. With suitable selections described with cuts on the kinematic variables, we realize statistical significance at about $5\sigma$ when the integrated luminosity is increased to about $100$ fb-1. We also perform a BDT analysis which results in such a large significance when the experimental analyses have lower luminosity as $\mathcal{L}\simeq 54$ fb-1. Such luminosity values can be reached on the first days of FCC-$h$ experiments.